OpenStudy (anonymous):

Determine the output voltage and curent of the transformer. Assume they are 100 efficient. 120 V 15 A primary coil 60 turns Secondary coil 20 turns

OpenStudy (anonymous):

The output voltage goes as the ratio of the turns. So, the secondary has one third as many turns, so the voltage is one third as much. Additionally, the power radiated away by the primary is the same as the power in the secondary. Since \[P = IV \] and V is decreased by a factor of three, I is increased by a factor of three.

OpenStudy (anonymous):

So, out put voltage is 1/3 then...output Voltage for 2nd is 40 V if power is the same, P=(120)(15) P=1800 1800=(I)(40 I=1800/40 I= 45 ??? right?

OpenStudy (anonymous):

Looks good to me

OpenStudy (anonymous):

Tnx Jem :) You probably helped me to pass my exam, whch is tomorrow Tnx!

OpenStudy (muhammad_nauman_umair):

As,Input voltage I1=15 Input voltage V1=120 N1=60 N2=20 N1/N2=V1/V2 so V2=40 ANd I1/I2=V2/V1 So I2=45amperes